The following present value factors are provided for use in this problem. Periods Present Value of $1 at 8% 0.9259 0.8573 0.7Multiple Choice Ο $2,702. Ο $1,820. Ο $39,302. Ο $(2,702). Ο $(1,820).
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=11600/1.08+11600/1.08^2+11600/1.08^3+11600/1.08^4+1200/1.08^4

=11600[1/1.08+1/1.08^2+1/1.08^3+1/1.08^4]+1200/1.08^4

=(11600*3.3121)+(1200*0.7350)

=39302.36

NPV=Present value of inflows-Present value of outflows

=39302.36-36600

=$2702(Approx).